# Kalman Filter Enhanced Group Relative Policy Optimization for Language Model Reasoning

### v4

@@ Abstract
The advantage function is a central concept in RL that helps reduce variance in policy gradient estimates. Recently, for language modeling, Group Relative Policy Optimization (GRPO) was proposed to compute the advantage for each output by subtracting the mean reward, as the baseline, for all outputs in the group. However, it can lead to high variance when the reward advantage is inaccurately estimated. In this work, we propose Kalman Filter Enhanced Group Relative Policy Optimization (KRPO) model, by using lightweight Kalman filtering to dynamically estimate the latent reward baseline and uncertainty. This filtering technique replaces the naive group mean, enabling more adaptive advantage normalization. Our method does not require additional learned parameters over GRPO. This approach offers a simple yet effective way to incorporate group-level uncertainty for advantage estimation, improving policy optimization in settings where highly dynamic reward signals are difficult to model for language models. Through the accuracies and rewards obtained from math question answering and reasoning, we show that using a more adaptive advantage estimation model, KRPO can improve the performance and show more stable return curves upon GRPO. The code is available at https://anonymous.4open.science/r/KRPO-E1D3.
Machine Learning, ICML
@@ 1 Introduction
To overcome this issue, Group Relative Policy Optimization (GRPO) (Shao et al., 2024) is proposed to simplify the prediction of reward baselines by taking the mean reward over a group of trajectories. However, this approach can lead to high variance if an inaccurate advantage is estimated when the reward is highly nonstationary. This may lead to unstable updates and slower convergence, especially in settings where the reward signal evolves rapidly with high variance over time.
To address this issue, we propose a dynamic advantage estimation method based on a one-dimensional Kalman filter, named as Kalman Filter Enhanced Group Relative Policy Optimization (KRPO). Instead of relying on the group mean, we treat the observed rewards as noisy measurements of a latent reward baseline signal and use the Kalman filter to model the underlying reward and noise trajectories. The resulting filtered baselines and uncertainties adapt over time, smoothing out noise while tracking underlying reward baseline values in the environment. This enables more accurate and robust advantage estimation with minimal computational overhead. Our method integrates naturally with the existing GRPO algorithm and does not require modification of policy networks. Provides a principled way to incorporate uncertainty into advantage estimation without introducing additional learning components or tuning objectives. We evaluate our approach by comparing it to the standard group-mean approach and show how it improves training stability. In summary, this paper makes the following contributions:
  • We introduce a lightweight Kalman filter approach for accurate advantage estimation. It can be seamlessly incorporated into the Group Relative Policy Optimization (GRPO) algorithm without introducing extra learning parameters.
  • We empirically demonstrate that the proposed KRPO algorithm can receive higher rewards than its standard group-mean counterparts across different datasets.
@@ 2.2 Reinforcement Learning to Improve Reasoning for LLMs
However, RLHF presents challenges, including the complexity of reward model training and the instability of reinforcement learning procedures. Alternative methods have been proposed to address these issues. Direct Preference Optimization (DPO) (Rafailov et al., 2023) eliminates the need for an explicit reward model by directly optimizing the policy to align with human preferences using a classification loss. This simplification leads to more stable and efficient training. Kahneman-Tversky Optimization (KTO) (Ethayarajh et al., 2024) incorporates principles from prospect theory to better model human decision-making biases. By aligning the training objective with human utility functions, KTO aims to improve the alignment of LLM outputs with human preferences. Reinforced Self-Training (ReST) (Gulcehre et al., 2023) adopts an offline reinforcement learning approach, generating training data from the model’s own outputs and refining the policy without the need for online interaction or extensive human feedback. Rank Responses with Human Feedback (RRHF) (Yuan et al., 2023) simplifies the alignment process by ranking multiple model-generated responses and optimizing the model to prefer higher-ranked outputs, reducing the dependency on complex reward models. Reinforcement Learning from AI Feedback (RLAIF) (Lee et al., 2023) replaces human feedback with evaluations from a separate LLM, enabling scalable alignment without the need for human-labeled data. Self-Play Fine-Tuning (SPIN) (Chen et al., 2024) leverages self-play mechanisms, allowing the model to generate and learn from its own data, progressively improving performance without additional human supervision. Group Relative Policy Optimization (GRPO) (Shao et al., 2024) introduces a group-based approach to policy optimization, where the model’s outputs are evaluated relative to a group baseline. Recent studies also have critically examined R1-Zero-style reinforcement learning (Liu et al., 2025).
These methods represent a focus towards more efficient and scalable approaches for aligning LLMs with human preferences, reducing dependence on extensive human feedback and complex reinforcement learning procedures. But there is limited exploration from the estimation of a more accurate advantage. Therefore, in this paper, we propose Kalman Filter Enhanced Group Relative Policy Optimization (KRPO), which replaces the group-mean-based advantage estimation in GRPO with a Kalman-filter-based one. By treating observed rewards as noisy measurements of a latent reward signal, the filter adaptively estimates a baseline that tracks the underlying uncertainties. With a non-parametric one-dimensional Kalman filter, we can have more accurate advantage estimation by filtering out the underlying reward baseline and uncertainty with minimal computational cost.
@@ 3.1 Problem Setting
We consider reinforcement learning in which an agent interacts with an environment modeled as a Markov Decision Process (MDP). In a timestep, the agent observes a state $s$, selects an action $a$ according to a policy $\pi(a\mid s)$, and receives a reward $r$. The goal is to maximize the expected cumulative return. Policy gradient methods rely on estimating the advantages to reduce the variance of gradient updates. The Group Relative Policy Optimization (GRPO) framework is applied to large language models by sampling multiple responses (i.e., action trajectories) from the policy given the same input prompt. A shared baseline is computed as the average return across all sampled responses for the same prompt, and the advantage for each output is calculated by subtracting the group sample mean. Although this reduces variance, the group mean may result in inaccurate estimates of action advantages, particularly when the rewards are very noisy. This can lead to unstable or suboptimal updates during training.
@@ 3.2 Kalman-Filtered Advantage Estimation
We propose replacing the naive group mean baseline with a Kalman filter that maintains an online estimate of the expected reward and its variance to obtain a more accurate estimate of the advantages. Each observed return $r_{i}$ is treated as a noisy measurement of the latent expected reward signal $\hat{x}_{i}$, which we cannot obtain directly. The subscript $i$ refers to the index of the i-th sampled observation.
The Kalman filter maintains a posterior estimate of the reward baseline $\hat{x}_{i|i}$ and its associated uncertainty $P_{i|i}$, both updated recursively. Here, the subscripts follow the standard Kalman notation: $\hat{x}_{i|i-1}$ denotes the prior estimate of “genuine” reward baseline $\hat{x_{i}}$ given observations up to step i1i-1, while $\hat{x}_{i|i}$ is the posterior estimate after incorporating the observation $r_{i}$. Similarly, $P_{i|i-1}$ and $P_{i|i}$ denote the corresponding uncertainty (variance) estimates before and after observing $r_{i}$.
@@
where $\varepsilon$ is a small constant for numerical computation stability. This formulation has two benefits: it adaptively centers the reward using a filtered baseline, and it normalizes the advantage by the estimated uncertainty, which helps stabilize the scale of policy updates and get a more reasonable advantage estimate.
@@ 3.3 KRPO Algorithm
Besides, each Kalman filter operation is applied to groups of rollouts. In the KRPO setting, although the reward values for individual samples are discrete, the grouped reward observations, which aggregate multiple rollouts, are not sparse and are much more continuous in distribution. According to the Central Limit Theorem (2), the sum of i.i.d. sampled rewards (with finite mean and finite variance) tends to follow a Gaussian distribution. This is consistent with the assumptions of our KRPO setting.
@@ 4 Connections of Group-mean and Kalman Filtering
The baseline computed in the GRPO model is the empirical mean of observed rewards. Specifically, for a group of returns $\{r_{i}\}_{i=1}^{n}$, the baseline is defined as
$\bar{r}=\frac{1}{n}\sum_{i=1}^{n}r_{i}.$ (7)
@@
This observation motivates the use of a Kalman filter, which accounts for both observation noise ($R>0$) and possible slight changes in the latent reward ($Q>0$), and dynamically adjusts the influence of new observations via the gain $K_{i}$. This leads to a more adaptive estimate of the reward advantage, making the Kalman filter a more general formulation compared to a simple average.
@@ 5.1 Implementation Details
Please note in the paper, we focus on math tasks, as math tasks are verifiable with well-defined answers. These tasks are very suitable for verify the effectiveness of KRPO also because they require the model to have adequate reasoning ability for resolving difficult math questions (even not easy for humans), and language model reasoning is the main focus of our KRPO paper. Reward function or reward model designing is not the main emphasis of our paper.
@@ 5.2.1 Overall Model Performance
Across both datasets and all difficulty levels, KRPO consistently outperforms baseline models. On the Arithmetic dataset, KRPO improves over GRPO by 2.296%, 5.361%, and 2.282% on Easy, Normal, and Hard subsets, respectively. Similarly, on the OpenMath-Instruct dataset, KRPO achieves substantial gains, with improvements of 4.623%, 12.373%, and 17.876% over GRPO. These improvements support that using a Kalman Filter provides a more stable and accurate advantage estimation (the only difference with GRPO), which benefits policy learning, especially on more challenging problems.
@@
We also show the statistical significance (p-value) of performance gaps between each model and the proposed KRPO model. All differences are statistically significant with p-values below the threshold 0.05, indicating that the observed improvements of KRPO to other models are large. These results collectively indicate that incorporating a Kalman Filter model for advantage estimation is especially beneficial in high-variance, complex reasoning environments. This aligns with the goal of improving the stability of the advantage computation during training. More performance comparisons with other models, including Direct Preference Optimization (DPO) (Rafailov et al., 2023) and fixed baseline models, can be found in the Appendix D.
@@ 5.2.2 Training Reward Curve Comparison between KRPO and GRPO
On both datasets and across all difficulties, the proposed KRPO algorithm can consistently converge faster and eventually reach much higher rewards. Interestingly, on the OpenMath-Instruct dataset, the performance gap between KRPO and GRPO becomes larger as problem difficulty increases. This trend suggests that KRPO provides more stable and accurate advantage estimates in harder reasoning scenarios, where the variance in reward signals is likely higher and a naive baseline, such as the mean reward, becomes less effective. This trend is also applied to the Arithmetic easy and normal subsets, but is not observed on the Arithmetic hard dataset. It is possibly caused by the difference in the arithmetic hard data distribution from other difficulty sets.
@@ 5.3 Analysis
In this sub-section, we analyze the effect of different hyper-parameters and observed phenomena (time efficiency analysis can be found in the Appendix G). For all experiments in the analysis section are conducted on Arithmetic data.
@@ 5.3.1 KL Loss Weights $\alpha$
Question {“type”: “Algebra”, “question”: “If $7^{4x}=343$, what is the value of $7^{4x-3}$”, “expected_answer”: “1”}
Model Thinking Process
@@ 5.3.2 Model Convergence against Process Noise and Measurement Noise
5.3.3 Case Study
We also provide a case study from OpenMath data as Tab. 2 to show the improved reasoning ability of the KRPO model over the GRPO baseline. From the case, we observe that GRPO can figure out the prior steps, e.g. “This implies $4x=3$, so $x=\frac{3}{4}$”, but it stops here and derives the wrong answer 343. In contrast, the proposed KRPO can get the correct answer 1. For this question, the KRPO can get 12 corrects out of 12 trials, while the GRPO can have 3 corrects out of 12. This is because KRPO can better award and punish correct and incorrect answers via a more accurate advantage estimation.
@@ 6 Conclusion
In this work, we improve the model performance over GRPO by introducing KRPO, a Kalman filter-based method for accurate advantage estimation. Our approach replaces the static mean with a Kalman-filtered baseline and uncertainty that adapt to latent reward dynamics. KRPO maintains the simplicity of GRPO and requires no modification to the policy network or additional learning parameters. Empirical results show that KRPO improves training convergence and policy performance across environments with dynamic and noisy rewards.
More broadly, KRPO demonstrates that principled uncertainty-aware estimation can be incorporated into policy optimization with minimal overhead, offering a practical alternative to estimate advantages. In future work, we plan to evaluate KRPO on tasks such as creative writing and summarization, where reward signals are inherently noisier and more subjective.
@@ Appendix A Datasets
Arithmetic Dataset (Open-Thought, 2024). It contains 100,000 arithmetic problems involving addition, subtraction, multiplication, and division. Each problem is represented as a natural language prompt and the target is the correct answer as a string. The arithmetic expressions include up to 6-digit numbers and up to 6 terms. We first split the 100,000 problems into a training set and a testing set using the standard 80%/20% split. During both training and testing, we filter out problems whose natural language prompt exceeds 128 characters. Then, we categorize the problems into three sub-datasets based on the number of digits and terms: the easy set includes problems with at most 3-digit numbers and at most 3 terms; the normal set includes problems with at most 5-digit numbers and at most 5 terms; and the hard set includes problems with at most 6-digit numbers and at most 6 terms.
OpenMath-Instruct Dataset (Toshniwal et al., 2024). This dataset contains 7,500 math problems from various topics, including Algebra, Geometry, Prealgebra, Precalculus, Number Theory, Intermediate Algebra, and Counting & Probability. Each problem is labeled with one of the five difficulty levels provided by the dataset. We split this dataset into a training set and a testing set using the same 80%/20% ratio. As with the first dataset, we filter out problems whose question text is longer than 128 characters during training and testing. We then group the problems into three sub-datasets based on difficulty level: the easy set includes problems with level $\leq$ 2; the normal set includes problems with level $\leq$ 3; and the hard set includes problems with level $\leq$ 5.
MATH500 (Lightman et al., 2023). MATH-500 is a test dataset of 500 math problems drawn from secondary/high school mathematics. Each instance consists of four fields: (1) problem, a mathematical question expressed in English with standard notation; (2) solution, a worked-out solution to the problem; (3) answer, the concise final answer; and (4) metadata including subject area (e.g. algebra, precalculus, geometry), difficulty level (integer from 1 to 5), and a unique identifier. Text fields (problem and solution) have average lengths in the thousands of characters; the answer field is very short. We also keep the standard 80%/20% split ratio for training and testing.
AIME (Veeraboina, 2024). AIME Problem Set (1983-2024) consists of 933 problems drawn from the American Invitational Mathematics Examination (AIME), covering the years 1983 through 2024. Each AIME contest contains 15 questions, and for each problem, the dataset provides the year, the problem number, the full problem statement, and the correct numerical answer. Problems are of high difficulty (designed for strong high school students), requiring non-trivial mathematical insight, covering a variety of topics such as algebra, number theory, combinatorics, geometry, etc. We divided the dataset into training and testing subsets with an 80%/20% split.

### v5

# Kalman Filter Enhanced GRPO for Reinforcement Learning-Based Language Model Reasoning

@@ Abstract
The advantage function is a central concept in RL that helps reduce variance in policy gradient estimates. For language modeling, Group Relative Policy Optimization (GRPO) was proposed to use the within-group sample mean as a baseline for advantage normalization. This estimator can be sensitive to small group size and rollout-level stochasticity, which may lead to suboptimal advantage estimates in some settings. In this paper, we propose Kalman Filter Enhanced Group Relative Policy Optimization (KRPO), a lightweight variant that treats per-group rewards as noisy observations of a latent prompt-level reward baseline and uses a 1D Kalman filter to estimate both the baseline and its uncertainty. KRPO introduces no additional learned parameters and can be integrated into GRPO with minimal computational overhead. On mathematical reasoning benchmarks, KRPO consistently improves training reward curves and final accuracy over GRPO. These results suggest that adaptive advantage estimation is a promising direction for critic-free reinforcement learning in language model reasoning. The code is available at https://github.com/billhhh/KRPO_LLMs_RL.
@@ 1 Introduction
To overcome this issue, Group Relative Policy Optimization (GRPO) (Shao et al., 2024) is proposed to simplify the prediction of reward baselines by taking the mean reward over a group of trajectories. However, GRPO uses the within-group sample mean as the baseline for advantage normalization. Although simple and effective, this estimator can be sensitive to small group sizes and rollout-level stochasticity, which can make advantage estimates unstable or less informative in some settings.
To address this limitation, we reinterpret group rewards as noisy observations of an underlying prompt-level reward signal and propose Kalman Filter Enhanced Group Relative Policy Optimization (KRPO), a lightweight 1D Kalman-filter-based baseline estimator that adaptively tracks both the latent baseline and its uncertainty. KRPO introduces no additional learned parameters and consistently improves performance and training stability over GRPO on mathematical reasoning benchmarks. These results suggest that adaptive baseline estimation is a promising direction for reinforcement learning-based language model reasoning. In summary, this paper makes the following contributions:
  • We introduce a lightweight Kalman filter approach for uncertainty-aware advantage estimation. It can be seamlessly incorporated into the Group Relative Policy Optimization (GRPO) algorithm without introducing extra learning parameters.
  • We empirically demonstrate that the proposed KRPO algorithm can receive higher rewards and accuracies than its standard group-mean counterparts across different mathematical reasoning benchmarks.
@@ 2.2 Reinforcement Learning to Improve Reasoning for LLMs
However, RLHF presents challenges, including the complexity of reward model training and the instability of reinforcement learning procedures. Alternative methods have been proposed to address these issues. Direct Preference Optimization (DPO) (Rafailov et al., 2023) eliminates the need for an explicit reward model by directly optimizing the policy to align with human preferences using a classification loss. This simplification leads to more stable and efficient training. Kahneman-Tversky Optimization (KTO) (Ethayarajh et al., 2024) incorporates principles from prospect theory to better model human decision-making biases. By aligning the training objective with human utility functions, KTO aims to improve the alignment of LLM outputs with human preferences. Reinforced Self-Training (ReST) (Gulcehre et al., 2023) adopts an offline reinforcement learning approach, generating training data from the model’s own outputs and refining the policy without the need for online interaction or extensive human feedback. Rank Responses with Human Feedback (RRHF) (Yuan et al., 2023) simplifies the alignment process by ranking multiple model-generated responses and optimizing the model to prefer higher-ranked outputs, reducing the dependency on complex reward models. Reinforcement Learning from AI Feedback (RLAIF) (Lee et al., 2023) replaces human feedback with evaluations from a separate LLM, enabling scalable alignment without the need for human-labeled data. Self-Play Fine-Tuning (SPIN) (Chen et al., 2024) leverages self-play mechanisms, allowing the model to generate and learn from its own data, progressively improving performance without additional human supervision. Group Relative Policy Optimization (GRPO) (Shao et al., 2024) introduces a group-based approach to policy optimization, where the model’s outputs are evaluated relative to a group baseline. Recent studies also have critically examined R1-Zero-style reinforcement learning (Liu et al., 2025). Dynamic sAmpling Policy Optimization (DAPO) (Yu et al., 2025) provides an open-source large-scale RL training system for LLM post-training.
These methods represent a focus towards more efficient and scalable approaches for aligning LLMs with human preferences, reducing dependence on extensive human feedback and complex reinforcement learning procedures. But there is limited exploration from the estimation of a more adaptive advantage. Therefore, in this paper, we propose Kalman Filter Enhanced Group Relative Policy Optimization (KRPO), which replaces the group-mean-based advantage estimation in GRPO with a Kalman-filter-based one. By treating observed rewards as noisy measurements of a latent reward signal, the filter adaptively estimates a baseline that tracks the underlying uncertainties. With a non-parametric one-dimensional Kalman filter, we can have more adaptive advantage estimation by filtering out the underlying reward baseline and uncertainty with minimal computational cost.
@@ 3.1 Problem Setting
We consider reinforcement learning in which an agent interacts with an environment modeled as a Markov Decision Process (MDP). In a timestep, the agent observes a state $s$, selects an action $a$ according to a policy $\pi(a\mid s)$, and receives a reward $r$. The goal is to maximize the expected cumulative return. Policy gradient methods rely on estimating the advantages to reduce the variance of gradient updates. The Group Relative Policy Optimization (GRPO) framework is applied to large language models by sampling multiple responses (i.e., action trajectories) from the policy given the same input prompt. A shared baseline is computed as the average return across all sampled responses for the same prompt, and the advantage for each output is calculated by subtracting the group sample mean. Although this reduces variance, when group size is limited and rollout outcomes are highly variable, this estimator may be unstable, which can in turn affect advantage normalization and policy updates.
@@ 3.2 Kalman-Filtered Advantage Estimation
We propose replacing the naive group mean baseline with a Kalman filter that maintains an online estimate of the expected reward and its variance to obtain a more adaptive estimate of the advantages. Each observed return $r_{i}$ is treated as a noisy measurement of the latent expected reward signal $\hat{x}_{i}$, which we cannot obtain directly. The subscript $i$ refers to the index of the i-th sampled observation.
The Kalman filter maintains a posterior estimate of the reward baseline $\hat{x}_{i|i}$ and its associated uncertainty $P_{i|i}$, both updated recursively. Here, the subscripts follow the standard Kalman notation: $\hat{x}_{i|i-1}$ denotes the prior estimate of latent prompt-level reward baseline $\hat{x_{i}}$ given observations up to step i1i-1, while $\hat{x}_{i|i}$ is the posterior estimate after incorporating the observation $r_{i}$. Similarly, $P_{i|i-1}$ and $P_{i|i}$ denote the corresponding uncertainty (variance) estimates before and after observing $r_{i}$.
@@
where $\varepsilon$ is a small constant for numerical computation stability. This formulation has two benefits: it adaptively centers the reward using a filtered baseline, and it normalizes the advantage by the estimated uncertainty, which helps stabilize the scale of policy updates and get a calibrated advantage scale.
@@ 3.3 KRPO Algorithm
Besides, each Kalman filter operation is applied to groups of rollouts. In the KRPO setting, although the reward values for individual samples are discrete, the grouped reward observations, which aggregate multiple rollouts, are not sparse. According to the Central Limit Theorem (2), the sum of i.i.d. sampled rewards (with finite mean and finite variance) tends to follow a Gaussian distribution. This is consistent with the assumptions of our KRPO setting.
@@ 4 Connections of Group-mean and Kalman Filtering
The baseline computed in the GRPO model is the empirical mean of observed rewards. Specifically, for a group of returns $\{r_{i}\}_{i=1}^{n}$, the baseline is defined as $\bar{r}=\frac{1}{n}\sum_{i=1}^{n}r_{i}$.
@@
The group mean can be viewed as a static equal-weight estimator, whereas KRPO uses a recursive uncertainty-aware estimator. Under simplifying assumptions, these two views are related. We use this connection as intuition for KRPO, rather than as a formal equivalence claim under arbitrary implementations.
@@ 5.1 Implementation Details
Scope of the empirical evaluation. Please note in the paper, we focus on math tasks, as math tasks are verifiable with well-defined answers. These tasks are suitable for verify the effectiveness of KRPO also because they require the model to have adequate reasoning ability for resolving difficult math questions. All main experiments are conducted with rule-based rewards in {0, 0.5, 1}. This setup is intentionally simple and isolates the effect of the baseline estimator from the confounding effects of learned reward models or subjective judges.
@@ 5.2.1 Overall Model Performance
Across both datasets and all difficulty levels, KRPO consistently outperforms baseline models. On the Arithmetic dataset, KRPO improves over GRPO by 2.296%, 5.361%, and 2.282% on Easy, Normal, and Hard subsets, respectively. Similarly, on the OpenMath-Instruct dataset, KRPO achieves substantial gains, with improvements of 4.623%, 12.373%, and 17.876% over GRPO. These improvements support that using a Kalman Filter provides a more adaptive and uncertainty-aware advantage estimation (the only difference with GRPO), which benefits policy learning, especially on challenging problems.
@@
We also show the statistical significance (p-value) of performance gaps between each model and the proposed KRPO model. All differences are statistically significant with p-values below the threshold 0.05. Paired comparisons on the test split favor KRPO over the reported baselines under the evaluation protocol used in this paper. These results indicate that KRPO is effective on the evaluated mathematical reasoning benchmarks and can be particularly helpful on harder tasks, where rollout variability is larger. This aligns with the goal of improving the stability of the advantage computation during training. More performance comparisons with other models, including Direct Preference Optimization (DPO) (Rafailov et al., 2023) and fixed baseline models, can be found in the Appendix D.
Because KRPO and GRPO differ only in the baseline / advantage estimator under otherwise matched settings, these gains suggest that adaptive, uncertainty-aware baseline estimation can be beneficial for policy learning on these tasks.
@@ 5.2.2 Training Reward Curve Comparison between KRPO and GRPO
On both datasets and across all difficulties, the proposed KRPO algorithm can consistently converge faster and eventually reach much higher rewards. Interestingly, on OpenMath-Instruct, the gap between KRPO and GRPO grows with task difficulty. One possible explanation is that harder problems induce greater rollout variability, making a raw group mean less reliable. This trend is also applied to the Arithmetic easy and normal subsets, but is not observed on the Arithmetic hard dataset. It can be caused by the difference in the arithmetic hard data distribution from other difficulty sets.
@@ 5.3 Analysis
In this sub-section, we analyze the effect of different hyper-parameters and observed phenomena. For all experiments in the analysis section are conducted on Arithmetic data.
5.3.1 Model Performance with Different Group Size and Rollout Stochasticity
(a) Model performance with different group sizes.
(b) Model performance with different random seeds.
We next examine when KRPO is most helpful. Figure 4(a) shows that the performance gap between KRPO and GRPO widens when fewer rollouts are available per prompt, suggesting that the raw group mean becomes less reliable in small-group regimes. Figure 4(b) shows that the gains are robust across random seeds. Together, these observations support a finite-sample interpretation of KRPO’s gains. For additional details, please refer to Appendix E and Appendix F.
@@ 6 Conclusion
In this work, we showed that replacing the raw group-mean baseline of GRPO with a lightweight Kalman-filter-based estimator can improve training stability and final performance on multiple mainstream mathematical reasoning tasks. KRPO preserves the simplicity of GRPO, without introducing additional learned parameters, and offers an uncertainty-aware alternative to mean-based advantage normalization.
More broadly, our results suggest that adaptive advantage estimation is a useful design direction for critic-free reinforcement learning for language model reasoning. Our current experiments do not fully characterize settings with learned reward models or generation tasks (such as creative writing, summarization or code writing). Extending KRPO to those directions is an important part for future work.
@@ Appendix A Datasets
Arithmetic Dataset [16]. It contains 100,000 arithmetic problems involving addition, subtraction, multiplication, and division. Each problem is represented as a natural language prompt and the target is the correct answer as a string. The arithmetic expressions include up to 6-digit numbers and up to 6 terms. We first split the 100,000 problems into a training set and a testing set using the standard 80%/20% split. During both training and testing, we filter out problems whose natural language prompt exceeds 128 characters. Then, we categorize the problems into three sub-datasets based on the number of digits and terms: the easy set includes problems with at most 3-digit numbers and at most 3 terms; the normal set includes problems with at most 5-digit numbers and at most 5 terms; and the hard set includes problems with at most 6-digit numbers and at most 6 terms.
OpenMath-Instruct Dataset [24]. This dataset contains 7,500 math problems from various topics, including Algebra, Geometry, Prealgebra, Precalculus, Number Theory, Intermediate Algebra, and Counting & Probability. Each problem is labeled with one of the five difficulty levels provided by the dataset. We split this dataset into a training set and a testing set using the same 80%/20% ratio. As with the first dataset, we filter out problems whose question text is longer than 128 characters during training and testing. We then group the problems into three sub-datasets based on difficulty level: the easy set includes problems with level $\leq$ 2; the normal set includes problems with level $\leq$ 3; and the hard set includes problems with level $\leq$ 5.
MATH500 [11]. MATH-500 is a test dataset of 500 math problems drawn from secondary/high school mathematics. Each instance consists of four fields: (1) problem, a mathematical question expressed in English with standard notation; (2) solution, a worked-out solution to the problem; (3) answer, the concise final answer; and (4) metadata including subject area (e.g. algebra, precalculus, geometry), difficulty level (integer from 1 to 5), and a unique identifier. Text fields (problem and solution) have average lengths in the thousands of characters; the answer field is very short. We also keep the standard 80%/20% split ratio for training and testing.
AIME [27]. AIME Problem Set (1983-2024) consists of 933 problems drawn from the American Invitational Mathematics Examination (AIME), covering the years 1983 through 2024. Each AIME contest contains 15 questions, and for each problem, the dataset provides the year, the problem number, the full problem statement, and the correct numerical answer. Problems are of high difficulty (designed for strong high school students), requiring non-trivial mathematical insight, covering a variety of topics such as algebra, number theory, combinatorics, geometry, etc. We divided the dataset into training and testing subsets with an 80%/20% split.